OpenStudy (anonymous):

The formula for degrees Celsius is c= 5/9 (f-32) where f stands for farenheight. What equation would you use to show the formula solved correctly for f?

OpenStudy (anonymous):

The answer choices are F= 5/9c+32 F= 9/5c +32 F= 5/9c -32 F= 9/5c -32

OpenStudy (anonymous):

Could you help @saifoo.khan or @lgbasallote ??

OpenStudy (saifoo.khan):

Sure @jrosesweet

OpenStudy (saifoo.khan):

So: c= 5/9 (f-32) We have to make "f" the subject.

OpenStudy (anonymous):

Alright...

OpenStudy (saifoo.khan):

In other words, we have to bring "f" on one side of the equation.

OpenStudy (saifoo.khan):

\[c= \frac59 (f-32) \] Multiply both sides with 9, \[9c = 5(f - 32)\]

OpenStudy (saifoo.khan):

Now, get rid of 5, by dividing 5 on both sides, \[\frac95c=f-32\]I'm pretty sure you can take from here?

OpenStudy (anonymous):

So it's f= 9/5c -32 ?? @saifoo.khan

OpenStudy (saifoo.khan):

Not exactly.

OpenStudy (saifoo.khan):

When we were in Kindergarden we studied that when we switch sides, the terms change their signs? Do you remember? :)

OpenStudy (anonymous):

Is it f= 9/5c + 32??

OpenStudy (saifoo.khan):

\[\Huge \color{yellow} \checkmark\]